A gun is fired vertically into a 1.40 kg block of wood at rest directly above it. If the bullet has a mass of 21.0 g and a speed of 210 m/s, how high will the block rise into the air after the bullet becomes embedded in it? 


1
Expert's answer
2021-10-18T11:10:25-0400
"1.4v=0.021(210)\\v=3.15\frac{m}{s}\\h=\frac{3.15^2}{2(9.8)}=0.506\ m"